# paper1.bib

### v1

@@ Abstract
Quasi-Newton methods form an important class of methods for solving nonlinear optimization problems. In such methods, first order information is used to approximate the second derivative. The aim is to mimic the fast convergence that can be guaranteed by Newton-based methods. In the best case, quasi-Newton methods will far outperform steepest descent and other first order methods, without the computational cost of calculating the exact second derivative. These convergence guarantees hold locally, which follows closely from the fact that if the objective function is strongly convex it can be approximated well by a quadratic function close to the solution. Understanding the performance of quasi-Newton methods on quadratic problems with a symmetric positive definite Hessian is therefore of vital importance. In the classic case, an approximation of the Hessian is updated at every iteration and exact line search is used. It is well known that the algorithm terminates finitely, even when the Hessian approximation is memoryless, i.e. requires only the most recent information. In this paper, we explore the possibilities in which reliance on exact line search and dependence on conjugate search directions can be relaxed, while preserving finite termination properties of quasi-Newton methods on quadratic problems.
@@ 1 Introduction
Given an initial point $x_{0}$, most methods to solve (QP) in a finite number of iterations generate mutually conjugate search directions $p_{k}$ with respect to $H$. At each iteration, the next iterate is obtained by taking the step
@@
In a quasi-Newton method, the search direction $p_{k}$ is found by solving the system $B_{k}p_{k}=-g_{k}$, where $g_{k}=g(x_{k})$ is the gradient at $x_{k}$ and $B_{k}\succ 0$ is some approximation of the Hessian.
The aim of this paper is to investigate the conditions under which a quasi-Newton method can achieve finite termination on (QP) without the need for exact line search. In Section 2 we briefly discuss quasi-Newton and Krylov subspace methods. In Section 3 we introduce the concept of a subspace Newton step. We generalize to a subspace quasi-Newton step in Section 4. Finally we present a quasi-Newton method that can achieve finite termination with the use of arbitrary step sizes in Section 5.
@@ 2 Background
In general, to solve (QP) we may find an exact solution in at most $n$ iterations with the use of conjugate search directions and exact line search. Assume that we have fixed some initial point $x_{0}\in\mathbb{R}^{n}$ and $\{p_{0},\dots,p_{n-1}\}$ are a set of conjugate vectors with respect to $H$, i.e. $p_{i}^{T}Hp_{j}=0$ for $i\neq j$. If exact line search is used to calculate the step size
$\alpha_{k}=-\frac{g_{k}^{T}p_{k}}{p_{k}^{T}Hp_{k}},$
@@
We will focus on Krylov subspace methods, where the search directions span a Krylov subspace as stated in the following definition.
@@
Suppose that we fix an initial point $x_{0}$ and search direction $p_{0}=-g_{0}$. Let $\hat{x}_{k}$ denote the minimizer of (QP) in the space $x_{0}+\mathcal{K}_{k}(g_{0},H)$ and $\hat{g}_{k}$ the gradient at that point. The differences $\hat{q}_{k}=\hat{x}_{k+1}-\hat{x}_{k}$ are mutually conjugate with respect to $H$ can be seen as search directions such that exact line search yields the unit step. Additionally $\mathcal{K}_{k}(g_{0},H)$ is spanned by the conjugate search directions $\{\hat{q}_{0},\dots,\hat{q}_{k-1}\}$, or equivalently by the gradients $\{\hat{g}_{0},\dots,\hat{g}_{k-1}\}$.
Once the initial point $x_{0}$ is fixed, a Krylov subspace method with exact line search follows a fixed path until $\hat{x}_{r}=x^{*}$. In the worst case $r=n$, however the algorithm will terminate at iteration $r$, where $r$ is the smallest integer such that $\mathcal{K}_{r}(g_{0},H)=\mathcal{K}_{r+1}(g_{0},H)$.
2.1 Generation of conjugate search directions
In order to solve (QP) using a Krylov subspace method, we must be able to generate search directions that are parallel to $\hat{q}_{k}$. We will denote parallel vectors by $p//p^{\prime}$. When exact line search is used on these parallel directions, the iterates are identical, and as such after $r$ iterations the algorithm will terminate.
@@ 2.1 Generation of conjugate search directions
$p_{0}=-g_{0}$
$p_{k}=-g_{k}+\frac{g_{k}^{T}Hp_{k-1}}{p_{k-1}^{T}Hp_{k-1}}p_{k-1}.$
For finite termination, it is necessary to use exact line search at each iteration to ensure that the gradients $g_{k}$ are orthogonal to all previous search directions so that the search directions are mutually conjugate.
The Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm is a quasi-Newton method, which in the case of a quadratic objective function, will also generate conjugate directions when exact line search is utilized.
Given an initial Hessian approximation $B_{0}\succ 0$, at each iteration a search direction is obtained by solving the system $B_{k}p_{k}=-g_{k}$ and the matrix is updated by
@@
The search directions obtained using CG, BFGS and memoryless BFGS all generate parallel search directions $p_{k}//\hat{q}_{k}$, and are therefore Krylov subspace methods.
2.2 Background results
The suggestion of a quasi-Newton methods was first made by Davidon, see [davidon] and then later further developed by Fletcher and Powell, see [fletcher]. For an overview on quasi-Newton methods see [nocedal, Chapter 4].
It has been shown that under certain conditions, $B_{k}$ will generate conjugate directions that are parallel, and with exact line search achieve finite termination in exactly $r$ iterations, see [forsgren2018exact, forsgren2015connection].
On finite termination of quasi-Newton methods, a result from Kolda, see [kolda], shows that finite termination can occur even when the matrix $B_{k}$ is not updated at every iteration if exact line search is used. A similar result from Powell, see [powell1, powell2], proves finite termination using unit steps as long as the update of $B_{k}$ is skipped in every other iteration.
@@ 3 Subspace Newton method
Suppose for some initial point $x_{0}$ we consider the $k$-th Krylov subspace $\mathcal{K}_{k}(g_{0},H)$. We may assume without loss of generality that $x_{0}$ is the origin. Then the Krylov space minimizer $\hat{x}_{k}$ will belong to the subspace $\mathcal{K}_{k}(g_{0},H)=\mathcal{K}_{k}(c,H)$.
For any arbitrary point $x\in\mathcal{K}_{k}(c,H)$ there exists a unique $(k-1)-th$ subspace Newton step $p^{N}_{k-1}(x)$ such that $x+p^{N}_{k-1}(x)=\hat{x}_{k}$.
Let a basis $\{s_{0},\dots,s_{k-1}\}$ of $\mathcal{K}_{k}(c,H)$ be chosen. If $g(x)=Hx+c$ is the gradient at the given point $x$, then the $(k-1)$-th subspace Newton step $p^{N}_{k-1}(x)$ can be expressed as
$p^{N}_{k-1}(x)=\left(\!\begin{array}[]{ccccc}s_{k-1}&s_{k-2}&\cdots&s_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right),$
for $\beta_{i}$, $i=0,\dots,k$, given by
$\left(\!\begin{array}[]{c}s_{k-1}^{T}\\ s_{k-2}^{T}\\ \vdots\\ s_{0}^{T}\end{array}\!\right)H\left(\!\begin{array}[]{ccccc}s_{k-1}&s_{k-2}&\cdots&s_{0}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k-1}\\ \beta_{k-2}\\ \vdots\\ \beta_{0}\end{array}\!\right)=-\left(\!\begin{array}[]{c}s_{k-1}^{T}\\ s_{k-2}^{T}\\ \vdots\\ s_{0}^{T}\end{array}\!\right)g(x).$
This follows from the fact that $\hat{x}_{k}$ is the unique point that minimizes (QP) over the Krylov subspace, or equivalently that the gradient at the minimizer $\hat{g}_{k}$ is orthogonal to $\mathcal{K}_{k}(c,H)$.
Suppose that instead of an arbitrary basis, we choose a conjugate basis with respect to $H$, $\{q_{0},\dots,q_{k-1}\}$, where $q_{i}//\hat{q}_{i}$. Taking into account that $q_{i}^{T}Hq_{j}=0$ for $i\neq j$, this yields
@@
Therefore it is simple to show that for this basis
$\beta_{i}=-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}$
for $i=0,\dots,k-1$. We call $\beta_{i}$ the Newton scaling of $q_{i}$ with respect to $x$. Indeed consider $\tilde{x}=x+\sum_{i=0}^{k-1}\beta_{i}q_{i}$. Then for any $j=0,\dots,k-1$
$g(\tilde{x})^{T}q_{j}=\left(g(x)+\sum_{i=0}^{k-1}\beta_{i}Hq_{i}\right)^{T}q_{j}=g(x)^{T}q_{j}+\beta_{j}q_{j}^{T}Hq_{j}=0.$
The $(k-1)$-th subspace Newton step may therefore be recovered using a Newton scaled conjugate basis for $\mathcal{K}_{k}(c,H)$.
@@
Let $x\in\mathcal{K}_{k}(c,H)$ and $p^{N}_{k-1}(x)$ the $k$-th subspace Newton step to $\hat{x}_{k}$. If $q_{k-1}//\hat{q}_{k-1}$ and $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, then the $k$-th subspace Newton step $p^{N}_{k}(x)$ is
$p^{N}_{k}(x)$ $=p^{N}_{k-1}(x)+\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}$
where
$\beta_{k}=-\frac{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}$
$\gamma_{k-1}=\frac{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}H\hat{g}_{k}}{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}.$
Proof. Since $\hat{x}_{k+1}\in\mathcal{K}_{k+1}(c,H)$, the $k$-th subspace Newton step $p^{N}_{k}(x)$ may be written as a linear combination of $\{\hat{g}_{k},q_{k-1},\dots,q_{0}\}$ where $q_{i}//\hat{q}_{i}$ for $i=0,\dots,k-1$. We assume without loss of generality that $q_{i},i=0,\dots,k-1$ has the Newton scaling with respect to $x$. Taking into account $\hat{g}_{k}^{T}Hq_{i}=0$, $i\leq k-2$, and $q_{i}^{T}\!Hq_{j}=0$, $i\neq j$, we obtain
@@
Then, $\beta_{i}=1$, $i=0,\dots,k-2$ by the Newton scaling.
Noting that $\hat{g}_{k}=g(x)+Hp^{N}_{k-1}(x)$, we may write $g(x)=\hat{g}_{k}-Hp^{N}_{k-1}(x)$. Since $p^{N}_{k-1}(x)=\sum_{i=0}^{k-1}q_{i}$ we have
$-\hat{g}_{k}^{T}g(x)$ $=$ $-\hat{g}_{k}^{T}(\hat{g}_{k}-Hp^{N}_{k-1}(x))=-\hat{g}_{k}^{T}\hat{g}_{k}+\hat{g}_{k}^{T}Hq_{k-1},$
$-q_{k-1}^{T}g(x)$ $=$ $-q_{k-1}^{T}(\hat{g}_{k}-Hp^{N}_{k-1}(x))=q_{k-1}^{T}Hq_{k-1},$
where the orthogonality of $\hat{g}_{k}$ to $q_{i}$, $i=0,\dots,k-1$ and $Hq_{i}$, $i=0,\dots,k-2$, as well as the conjugacy of the $q_{i}$s have been used.
@@
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \beta_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}+\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}Hq_{k-1}\end{array}\!\right).$
Let $\beta_{k-1}=1+\gamma_{k-1}$, then this is reduced to
@@
The $k$-th subspace Newton step to $\hat{x}_{k+1}$ is therefore formed by
$p^{N}_{k}(x)=\sum_{i=0}^{k-1}q_{i}+\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}=p^{N}_{k-1}(x)+\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}$
where $\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}=\hat{q}_{k}$.
Note that $\beta_{k}$ and $\gamma_{k-1}q_{k-1}$ are independent of the scaling of $q_{k-1}$, thus the formula for the $k$-th subspace Newton step is valid for any $q_{k-1}//\hat{q}_{k-1}$.

### v2

@@ Abstract
Quasi-Newton methods form an important class of methods for solving nonlinear optimization problems. In such methods, first order information is used to approximate the second derivative. The aim is to mimic the fast convergence that can be guaranteed by Newton-based methods. In the best case, quasi-Newton methods will far outperform steepest descent and other first order methods, without the computational cost of calculating the exact second derivative. These convergence guarantees hold locally, which follows closely from the fact that, if the objective function is strongly convex, it can be approximated well by a quadratic function close to the solution. Understanding the performance of quasi-Newton methods on quadratic problems with a symmetric positive definite Hessian is therefore of vital importance. In the classic case, an approximation of the Hessian is updated at every iteration and exact line search is used. It is well known that the algorithm terminates finitely, even when the Hessian approximation is memoryless, i.e. requires only the most recent information. In this paper, we explore the possibilities in which reliance on exact line search and dependence on conjugate search directions can be relaxed, while preserving finite termination properties of quasi-Newton methods on quadratic problems. We show that it suffices to create a memoryless quasi-Newton matrix based on two vectors to give ability to compute a Newton direction within a finite number of iterations, independent of step lengths. It is unnecessary for the quasi-Newton approximation to act as the Hessian on the full space.
@@ 1 Introduction
Given an initial point $x_{0}$, (QP) may be solved in a finite number of iterations by an iterative method that generates mutually conjugate search directions $p_{k}$ with respect to $H$. At each iteration, the next iterate is obtained by taking the step
@@
In a quasi-Newton method, the search direction $p_{k}$ is found by solving the system $B_{k}p_{k}=-g_{k}$, where $g_{k}=g(x_{k})$ is the gradient at $x_{k}$ and $B_{k}=B_{k}^{T}\succ 0$ is some approximation of the Hessian $H$.
The aim of this paper is to investigate the conditions under which a quasi-Newton method can achieve finite termination on (QP) without the need for exact line search. In Section 2 we briefly discuss Krylov subspaces and Krylov subspace methods such as the method of conjugate gradients as well as quasi-Newton methods. In Section 3 we introduce the concept of a subspace Newton step within a given Krylov subspace as well as an expanding Krylov subspace. We generalize to a subspace quasi-Newton step in Section 4, showing that it suffices to create a quasi-Newton matrix based on at most two vectors. Finally we present an iterative memoryless quasi-Newton method based on at most two vectors, in addition to the gradient, that will compute the Newton step in a finite number of iterations with the use of arbitrary step sizes in Section 5.
@@ 2 Background
In general, to solve (QP) we may find an exact solution in at most $n$ iterations with the use of conjugate search directions and exact line search. Given a point $x\in\mathbb{R}^{n}$ and a search direction $p$, exact line search can be used to calculate the optimal step length
$\alpha=-\frac{g(x)^{T}p}{p^{T}Hp}.$
@@
The focus of this paper will be on Krylov subspace methods, where the search directions span a Krylov subspace as stated in the following definition.
@@
Suppose that we fix an initial point $x_{0}$, search direction $p_{0}=-g_{0}$ and consider the Krylov subspaces $\mathcal{K}_{k}(g_{0},H)$. Any Krylov subspace method will follow a fixed path to minimizers over expanding affine spaces until termination occurs in a finite number of iterations.
Definition 2.2
Given $x_{0}\in\mathbb{R}^{n}$, let $r$ be the smallest integer such that $\mathcal{K}_{r}(g_{0},H)=\mathcal{K}_{r+1}(g_{0},H)$. For $k\leq r$, denote the minimizer of (QP) over the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$ as $\hat{x}_{k}$. The gradient at the minimizer is denoted by $\hat{g}_{k}$ and the search direction generated by any Krylov subspace method is parallel to the difference $\hat{q}_{k-1}=\hat{x}_{k}-\hat{x}_{k-1}$.
The search directions $\hat{q}_{0},\dots,\hat{q}_{k-1}$ are mutually conjugate with respect to $H$ and form a basis for $\mathcal{K}_{k}(g_{0},H)$. Additionally, the gradients $\hat{g}_{0},\dots,\hat{g}_{k-1}$ form an orthogonal basis for the same subspace. For $r$ as defined above, the minimizer $\hat{x}_{r}$ is exactly the global minimizer $x^{*}$ of (QP). For a more comprehensive overview of Krylov subspace methods, see, e.g [nocedal, Chapter 5], [Saa95, Chapter 6].
2.1 Generation of conjugate search directions using exact line search
In order to solve (QP) using a Krylov subspace method, search directions that are parallel to $\hat{q}_{k}$ must be generated. Parallel vectors will be denoted by $p//p^{\prime}$. When exact line search is used on these parallel directions, the iterates are identical, and as such after $r$ iterations the algorithm will terminate.
@@ 2.1 Generation of conjugate search directions using exact line search
$p_{0}=-g_{0},\quad p_{k}=-g_{k}+\frac{g_{k}^{T}Hp_{k-1}}{p_{k-1}^{T}Hp_{k-1}}p_{k-1}=-g_{k}+\frac{g_{k}^{T}g_{k}}{g_{k-1}^{T}g_{k-1}}p_{k-1}.$
For finite termination, it is necessary to use exact line search at each iteration to ensure that the gradients $g_{k}$ are orthogonal to all previous search directions so that the search directions are mutually conjugate. For a detailed description of the methods of conjugate gradients, see, e.g. Saad [Saa95, Chapter 6.7].
2.2 Quasi-Newton methods
The Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm is a quasi-Newton method, which in the case of a quadratic objective function, will also generate conjugate directions when exact line search is utilized. Given an initial Hessian approximation $B_{0}\succ 0$, at each iteration a search direction is obtained by solving the system $B_{k}p_{k}=-g_{k}$ and the matrix is updated by
@@ 2.2 Quasi-Newton methods
The different methods CG, BFGS and memoryless BFGS all generate parallel search directions $p_{k}//\hat{q}_{k}$ if exact line search is used, and are therefore Krylov subspace methods [huang].
The suggestion of a quasi-Newton methods was first made by Davidon [davidon] and then later further developed by Fletcher and Powell [fletcher]. For an overview on quasi-Newton methods, see, e.g., [nocedal, Chapter 4], [GMW19, Chapter 4.5] and [Fle81, Chapter 3.5]. There is a vast literature on quasi-Newton methods, which is beyond the scope of the current manuscript. We will point to the BFGS method and the memoryless BFGS method, which are of relevance for our results.
It has been shown that under certain conditions, a quasi-Newton matrix $B_{k}$ will generate conjugate directions that are parallel, and with exact line search achieve finite termination in exactly $r$ iterations [forsgren2018exact, forsgren2015connection].
On finite termination of quasi-Newton methods, a result from Kolda, O’Leary and Nazareth [kolda], shows that finite termination can occur even when the quasi-Newton matrix $B_{k}$ is not updated at every iteration if exact line search is used. A similar result from Powell [powell1, powell2], proves finite termination using unit steps as long as the update of the quasi-Newton matrix $B_{k}$ is skipped in every other iteration.
@@ 3 Subspace Newton method
Suppose for some initial point $x_{0}$ we consider the Krylov subspace $\mathcal{K}_{k}(g_{0},H)$. Then the Krylov space minimizer $\hat{x}_{k}$ will belong to the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$. For any arbitrary point $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ there exists a unique subspace Newton step $p^{N}_{k-1}(x)\in\mathcal{K}_{k}(g_{0},H)$ such that $x+p^{N}_{k-1}(x)=\hat{x}_{k}$. Suppose $S=\{s_{0},\dots,s_{k-1}\}$ has columns that form a basis of $\mathcal{K}_{k}(g_{0},H)$. If $g(x)=Hx+c$ is the gradient at the given point $x$, then the subspace Newton step $p^{N}_{k-1}(x)$ can be expressed as
$p^{N}_{k-1}(x)=S\beta,$
for $\beta$ given by
$S^{T}HS\beta=-S^{T}g(x).$
This follows from the fact that $\hat{x}_{k}$ is the unique point that minimizes (QP) over the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$, or equivalently that the gradient at the minimizer $\hat{g}_{k}$ is orthogonal to $\mathcal{K}_{k}(g_{0},H)$. The subspace Newton step will be fundamental in our analysis in that the relative change of the directional derivatives of the gradient along the subspace Newton step with respect to all conjugate directions can be characterized by one steplength $\alpha$, as stated in the following lemma.
Lemma 3.1
For $k\leq r$, let $q_{0},\dots,q_{k-1}$ be a basis for $\mathcal{K}_{k}(g_{0},H)$ where $q_{i}//\hat{q}_{i}$. For any $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, the subspace Newton step $p^{N}_{k-1}(x)$ to $\hat{x}_{k}$ is
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}.$
Additionally, for any $\alpha\in\mathbb{R}$
$g(x+\alpha p_{k-1}^{N}(x))^{T}q_{j}=(1-\alpha)g(x)^{T}q_{j},\quad j=0,\dots,k-1.$
Proof. Since $q_{i}//\hat{q}_{i},i=0,\dots,k-1$, the basis $q_{0},\dots,q_{k-1}$ is mutually conjugate with respect to $H$. We may therefore represent the subspace Newton step as $p^{N}_{k-1}(x)=\sum_{i=0}^{k-1}\beta_{i}q_{i}$ where $\beta_{i},i=0,\dots,k-1$ are the solutions to the system
@@
Therefore it is straightforward to show that
$\beta_{i}=-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}},\qquad i=0,\dots,k-1$
and the subspace Newton step is
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}.$
Additionally, for any $\alpha\in\mathbb{R}$ and $j=0,\dots,k-1$,
$g(x+\alpha p_{k-1}^{N}(x))^{T}\!q_{j}$ $=(g(x)-\alpha\sum_{i=0}^{k-1}\frac{g(x)^{T}\!q_{i}}{q_{i}^{T}Hq_{i}}Hq_{i})^{T}q_{j}$
$=g(x)^{T}\!q_{j}-\alpha\frac{g(x)^{T}\!q_{j}}{q_{j}^{T}Hq_{j}}q_{j}^{T}Hq_{j}=(1-\alpha)g(x)^{T}q_{j},$
where the conjugacy of the vectors $q_{0},\dots,q_{k-1}$ has been used.
For a point $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, the subspace Newton step to $\hat{x}_{k}$ may therefore be recovered using a Newton scaled conjugate basis for $\mathcal{K}_{k}(g_{0},H)$. Once the step $p^{N}_{k-1}(x)$ is known explicitly, it is possible to recover $\hat{x}_{k}$ and therefore the gradient $\hat{g}_{k}$. Since $\hat{g}_{k}$ is orthogonal to $\mathcal{K}_{k}(g_{0},H)$, the vectors $\{\hat{g}_{k},q_{k-1},\dots,q_{0}\}$ form a basis for $\mathcal{K}_{k+1}(g_{0},H)$. Using such a basis, we may construct the subspace Newton step $p^{N}_{k}(x)$ to the minimizer $\hat{x}_{k+1}\in x_{0}+\mathcal{K}_{k+1}(g_{0},H)$ for any $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$.
@@
For $k<r$, let $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ and let $p^{N}_{k-1}(x)$ be the subspace Newton step to $\hat{x}_{k}$. If $q_{k-1}//\hat{q}_{k-1}$ and $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, then the subspace Newton step $p^{N}_{k}(x)$ to $\hat{x}_{k+1}$ is
$p^{N}_{k}(x)$ $=\hat{q}_{k}+p_{k-1}^{N}(x)$
with
$\hat{q}_{k}=\frac{1}{\hat{\sigma}_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)\text{and }\hat{\sigma}_{k}=\frac{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}}.$
Proof. Since $\hat{x}_{k+1}\in x_{0}+\mathcal{K}_{k+1}(g_{0},H)$, the subspace Newton step $p^{N}_{k}(x)$ may be written as a linear combination of $\hat{g}_{k},q_{k-1},\dots,q_{0}$ where $q_{i}//\hat{q}_{i}$ for $i=0,\dots,k-1$ which forms a basis for $\mathcal{K}_{k+1}(g_{0},H)$. Taking into account that $\hat{g}_{k}^{T}Hq_{i}=0$, $i\leq k-2$, and $q_{i}^{T}\!Hq_{j}=0$, $i\neq j$, yields
@@
Then,
$\beta_{i}=-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}},\qquad i=0,\dots,k-2$
is the Newton scaling with respect to $x$.
Noting that $\hat{g}_{k}=g(x)+Hp^{N}_{k-1}(x)$, we may write $g(x)=\hat{g}_{k}-Hp^{N}_{k-1}(x)$. Since
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}.$
by Lemma 3.1, we have
$-\hat{g}_{k}^{T}g(x)=-\hat{g}_{k}^{T}(\hat{g}_{k}-Hp^{N}_{k-1}(x))=-\hat{g}_{k}^{T}\hat{g}_{k}-\frac{g(x)^{T}q_{k-1}}{q_{k-1}^{T}Hq_{k-1}}\hat{g}_{k}^{T}Hq_{k-1},$
where the orthogonality of $\hat{g}_{k}$ to $Hq_{i},i=0,\dots,k-2$ is used.
@@
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}H\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\beta_{k}\\ \beta_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}-\frac{g(x)^{T}q_{k-1}}{q_{k-1}^{T}Hq_{k-1}}\hat{g}_{k}^{T}Hq_{k-1}\\ -q_{k-1}^{T}g(x)\end{array}\!\right).$
Let $\beta_{k-1}=-\frac{g(x)^{T}q_{k-1}}{q_{k-1}^{T}Hq_{k-1}}+\gamma_{k-1}$, then this is reduced to
@@
The subspace Newton step to $\hat{x}_{k+1}$ is therefore formed by
$p^{N}_{k}(x)=\beta_{k}\hat{g}_{k}+\gamma_{k-1}q_{k-1}+\sum_{i=0}^{k-1}-\frac{g(x)^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}=\frac{1}{\hat{\sigma}_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)+p_{k-1}^{N}(x)$
where
$\hat{\sigma}_{k}=-\frac{1}{\beta_{k}}=\frac{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}}.$